### v1

# Local MAP Sampling for Diffusion Models

Abstract
Diffusion Posterior Sampling (DPS) provides a principled Bayesian approach to inverse problems by sampling from $p(x_{0}\mid y)$. However, in practice, the goal of inverse problem solving is not to cover the posterior but to recover the most accurate reconstruction, where optimization-based diffusion solvers often excel despite lacking a clear probabilistic foundation. We introduce Local MAP Sampling (LMAPS), a new inference framework that iteratively solving local MAP subproblems along the diffusion trajectory. This perspective clarifies their connection to global MAP estimation and DPS, offering a unified probabilistic interpretation for optimization-based methods. Building on this foundation, we develop practical algorithms with a probabilistically interpretable covariance approximation, a reformulated objective for stability and interpretability, and a gradient approximation for non-differentiable operators. Across a broad set of image restoration and scientific tasks, LMAPS achieves state-of-the-art performance, including $\geq 2$ dB gains on motion deblurring, JPEG restoration, and quantization, and >1.5>1.5 dB improvements on inverse scattering benchmarks.
1 Introduction
[…]
However, we argue that posterior sampling is not fully aligned with the objectives of inverse problem solving. Evaluation protocols for inverse problems reflect this philosophy: they typically compare reconstructions against a single ground-truth reference and provide no credit for coverage or diversity. For example, in medical imaging, the primary objective is to obtain the most accurate reconstruction or best estimate of the underlying image, rather than to generate diverse samples from the posterior distribution.
Optimization-based approaches—such as Resample (song2023solving, ), DiffPIR (zhu2023denoising, ), DCDP (li2024decoupled, ), and DMPlug (wang2024dmplug, )—have shown strong performance by alternating between denoising, optimization, and resampling to address inverse problems. Unlike DPS, which attempts to sample from the posterior distribution $p(x_{0}\mid y)$, optimization-based approaches prioritize reconstruction performance over distributional faithfulness. Nevertheless, their probabilistic foundation remains unclear: if the iterative procedure converges to the global MAP solution, i.e., $\arg\max p(x_{0}\mid y)$, would it still be consistent with DPS? Clarifying this foundation could provide both a principled interpretation and a stronger theoretical basis for optimization-based methods.
In this work, we argue that the optimization steps in these methods inherently solve a local MAP problem. But the resulting solutions neither converge to the global MAP nor equivalent to posterior sampling. Instead, they are more likely to reflect a trade-off between the two.
Our main contributions are summarized as follows:
  • Theoretical. We formulate Local MAP Sampling (LMAPS), a new inference framework that iteratively solves local maximum-a-posteriori subproblems along the diffusion trajectory. We analyze its relationship to global MAP and DPS, and show that LMAPS unifies Tweedie Moment Projected Diffusion (TMPD) and optimization-based inverse problem methods under a single framework. The relationship between LMAPS and existing methods are presented in Figure 1.
  • Methodological. To address inverse problems, we propose a probabilistically interpretable covariance approximation, replacing heuristic choices in existing solvers. We further introduce an objective reformulation for interpretable parameters and improved stability, and develop a gradient approximation strategy for non-differentiable operators (e.g., JPEG restoration, quantization).
  • Empirical. LMAPS is validated on 10 image restoration tasks (linear, nonlinear, non-differentiable) and 3 scientific inverse problems. It achieves the best results in $46/60$ FFHQ/ImageNet cases with $\geq 2$ dB PSNR gains on motion deblurring, JPEG restoration, and quantization, while being more efficient than DAPS. On scientific tasks, LMAPS consistently attains the highest PSNR, including >1.5>1.5 dB gains on 3 linear inverse scattering tasks.
Figure 1: Comparison of LMAPS with other methods. (a). The relationship between different alignment approaches; (b). The generation process of unconditional diffusion model; (c). The generation process of LMAPS.
2 Background
Unconditional diffusion models. The goal of diffusion model is to sample from an unknown distribution $\pi_{0}(x_{0})$ given a training dataset $\mathcal{D}=\{x_{0}^{i}\}_{i=1}^{N}$. Given a data point $x_{0}\sim\pi_{0}$ and a time step $t$, a noisy datapoint is sampled from the transition kernel: $p_{t}(x_{t}\mid x_{0})=\mathcal{N}(x_{t};\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})$. Diffusion process is built by mixture of densities: $p_{t}(x_{t})=\int p_{t}(x_{t}\mid x_{0})\pi_{0}(x_{0})dx_{0}$, and DDIM samples $\pi_{0}(x_{0})$ by running an iterative process $p_{t}(x_{t})$ from time $t=T$ to $t=0$ with the initial condition $x_{T}\sim p(x_{T})$:
$x_{t-\Delta t}=g(m_{0\mid t}(x_{t}),x_{t},\epsilon),\quad\epsilon\sim\mathcal{N}(0,\mathbb{I})$ (1)
where $\epsilon\sim\mathcal{N}(0,\mathbb{I})$ is the fresh noise added at the inference time, $m_{0\mid t}(t,x)=\mathbb{E}[x_{0}\mid x_{t}]$ is the ideal denoiser, and we define:
$g(\xi,x_{t},\epsilon):=\alpha_{t-\Delta t}\xi+\sigma_{t-\Delta t}(\sqrt{1-\rho_{t}^{2}}\frac{x_{t}-\alpha_{t}\xi}{\sigma_{t}}+\rho_{t}\epsilon),$ (2)
The goal of posterior sampling is to generate samples under some condition $y$, i.e., sample $x_{0}$ from a posterior distribution, $\pi_{0\mid y}(x_{0}\mid y)$, where $y$ could be class labels, measurements or text information, for example. There are two main approaches to do posterior sampling with diffusion prior: diffusion posterior sampling and Decoupled Annealing Posterior Sampling (DAPS).
Diffusion Posterior Sampling (DPS). DPS generate $x_{0}\sim\pi_{0\mid y}(x_{0}\mid y)$ by running an iterative process $p_{t\mid y}(x_{t}\mid y)$ from time $t=T$ to $t=0$ with the initial condition $x_{T}\sim p(x_{T}\mid y)$:
$x_{t-\Delta t}=g(m_{0\mid t,y}(t,x_{t},y),x_{t},\epsilon),\quad\epsilon\sim\mathcal{N}(0,\mathbb{I}),$ (3)
where $m_{0\mid t,y}(t,x_{t},y)=\mathbb{E}[x_{0}\mid x_{t},y]$ is the conditional denoiser. According Tweedie’s formula,
$\mathbb{E}[x_{0}\mid x_{t},y]=m_{0\mid t}+\frac{\sigma_{t}^{2}}{\alpha_{t}}\nabla_{x_{t}}\log p(y\mid x_{t}).$ (4)
Eq. 4 connects the conditional denoiser $\mathbb{E}[x_{0}\mid x_{t},y]$ with the unconditional denoiser $\mathbb{E}[x_{0}\mid x_{t}]$. However, the additional term $\nabla_{x_{t}}\log p(y\mid x_{t})$ is still intractable. One can train a neural network to approximate $\nabla_{x_{t}}\log p(y\mid x_{t})$, like classifier guidance (dhariwal2021diffusion, ). Training-free guidance, such as in (chung2022diffusion, ), usually approximates $\nabla_{x_{t}}\log p(y\mid x_{t})$ by a convenient single-sample approximation, $p(y\mid x_{t})\approx p(y\mid m_{0\mid t}(x_{t}))$, according to chain rule:
$\nabla_{x_{t}}\log p(y\mid x_{t})\approx\nabla_{x_{t}}m_{0\mid t}(t,x_{t})\nabla_{m_{0\mid t}}\log p(y\mid m_{0\mid t}(t,x_{t})).$ (5)
Decoupled Annealing Posterior Sampling (DAPS) (zhang2025improving, ). Alternatively, DAPS developed a new framework to sample $x_{0}\sim\pi_{0\mid y}(x_{0}\mid y)$, which is given by the following iterations:
$x_{0\mid t,y}$ $\sim p(x_{0}\mid x_{t},y)$ (6)
$x_{t-\Delta t}$ $=g(x_{0\mid t,y},x_{t},\epsilon),\quad\epsilon\sim\mathcal{N}(0,\mathbb{I}).$
Approximate posterior samples $x_{0\mid t,y}$ are obtained at each diffusion step using Langevin dynamics.
Algorithm 1 DPS
1:Input: $x_{t_{N}}\sim\pi_{T}$
2:for $k=N$ to $1$ do
3:  $\tilde{x}_{0}=\mathbb{E}[x_{0}\mid x_{t_{k}},y]$
4:  $\epsilon\sim\mathcal{N}(0,\mathbb{I})$
5:  $x_{{}_{t_{k-1}}}=g(\tilde{x}_{0},x_{t_{k}},\epsilon)$
6:end for
7:return $x_{0}$
Algorithm 2 DAPS
1:Input: $x_{t_{N}}\sim\pi_{T}$
2:for $k=N$ to $1$ do
3:  $\tilde{x}_{0}\sim p(x_{0}\mid x_{t_{k}},y)$
4:  $\epsilon\sim\mathcal{N}(0,\mathbb{I})$
5:  $x_{t_{k-1}}=g(\tilde{x}_{0},x_{t_{k}},\epsilon)$
6:end for
7:return $x_{0}$
Algorithm 3 LMAPS
1:Input: $x_{t_{N}}\sim\pi_{T}$
2:for $k=N$ to $1$ do
3:  $\tilde{x}_{0}=\arg\max p(x_{0}\mid x_{t_{k}},y)$
4:  $\epsilon\sim\mathcal{N}(0,\mathbb{I})$
5:  $x_{t_{k-1}}=g(\tilde{x}_{0},x_{t_{k}},\epsilon)$
6:end for
7:return $x_{0}$
Figure 2: Comparison of inference algorithm between DPS, DAPS and LMAPS.
3 Local MAP Sampling
3.1 Local MAP and global MAP
Global MAP. In Bayesian inference, the maximum a posteriori (MAP) estimate is defined as the single configuration that maximizes the posterior probability,
$x_{0}^{\text{MAP}}:=\arg\max_{x_{0}}p(x_{0}\mid y).$ (7)
We refer to this as the global MAP, since it directly targets the mode of the full posterior distribution after conditioning on the observation $y$. Unlike posterior sampling methods (e.g., DPS or DAPS), which produce diverse draws from $p(x_{0}\mid y)$, global MAP produces a single point estimate. This estimate prioritizes fidelity and certainty over diversity, offering a principled way to recover a solution that best aligns with both the diffusion prior and the measurement model.
Local MAP. Directly solving for $x_{0}^{\text{MAP}}$ in high-dimensional, non-convex posteriors can be computationally intractable. Instead, we consider a sequence of local MAP problems, which implemented by DDIM-like iteration from time $t=T$ to $t=0$ with the initial condition $x_{T}\sim p(x_{T}\mid y)$:
$x_{0}^{*}(t,x_{t},y)$ $:=\operatorname*{arg\,max}p(x_{0}\mid x_{t},y),$ (8a)
$x_{t-\Delta t}$ $=g(x_{0}^{*},x_{t},\epsilon),\quad\epsilon\sim\mathcal{N}(0,\mathbb{I}).$ (8b)
Eq. 8a and Eq. 8b correspond to the local MAP step and the DDIM update step, respectively. In particular, the local MAP step is equivalent to:
$x_{0}^{*}(t,x_{t},y)=\arg\min\{-\log p(x_{0}\mid x_{t})-\log p(y\mid x_{0})\}.$ (9)
This optimization problem can be solved via gradient descent if $\log p(x_{0}\mid x_{t})$ and $\log p(y\mid x_{0})$ are known and differentiable, although in practice we approximate $p(x_{0}\mid x_{t})$ as discussed in Sec. 4.
Figure 3: Comparison of LMAPS, DPS, DAPS and Global MAP on 2D synthetic data. LMAPS is less likely to generate samples in the between-mode regions or low-density regions.
3.2 The difference between DPS, local MAP and global MAP
One might expect that the iteration in Eq. 8 can be used to sample from the posterior $p(x_{0}\mid y)$ or converge to global MAP $\arg\max p(x_{0}\mid y)$. Unfortunately, this is generally not the case.
DPS vs. local MAP. DPS evolves $x_{t}$ by using the conditional mean $m_{0\mid t,y}(t,x_{t},y)=\mathbb{E}[x_{0}\mid x_{t},y]$ inside the DDIM update (Eq. 3), whereas local MAP replaces the mean with the conditional mode: $x_{0}^{*}(t,x_{t},y)=\arg\max p(x_{0}\mid x_{t},y)$, and then plugs $x_{0}^{*}$ into the same $g(\cdot)$ transition (Eq. 8). Consequently, replacing $\mathbb{E}[x_{0}\mid x_{t},y]$ with $\arg\max p(x_{0}\mid x_{t},y)$ alters the forward operator acting on $p_{t\mid y}(x_{t})$ and does not preserve the posterior marginals $p_{t\mid y}$.
When are DPS and local MAP equivalent? These two coincide if and only if $\mathbb{E}[x_{0}\mid x_{t},y]=\arg\max p(x_{0}\mid x_{t},y)$, for example if $p(x_{0}\mid x_{t},y)$ is (uni-variate or multi-variate) Gaussian. The condition holds, e.g., in linear-Gaussian inverse problems with a Gaussian diffusion prior approximation (quadratic negative log-density), with detailed discussion in Sec. 4. Outside of this setting (nonlinear forward models, heavy-tailed likelihoods, mixture-like priors), the posterior $p(x_{0}\mid x_{t},y)$ is non-Gaussian and the two updates generally differ. With non-Gaussian $p(x_{0}\mid x_{t},y)$, local MAP introduces a mode-seeking bias and does not reproduce posterior sampling.
Local MAP vs. global MAP. Global MAP is the single maximizer $x_{0}^{\text{MAP}}=\arg\max p(x_{0}\mid y)$. Local MAP instead solves, at each time $t$, a conditioned optimization (Eq. 9): $x_{0}^{*}(t,x_{t},y)=\arg\max p(x_{0}\mid x_{t},y)$. Because $x_{t}$ itself depends on the entire past trajectory (initialization, noise schedule, and random seeds), the sequence of local maximizers need not approach the global maximizer of $p(x_{0}\mid y)$ as $t\downarrow 0$.
In summary, DPS targets $p(x_{0}\mid y)$, and LMAPS targets $\arg\max p(x_{0}\mid x_{t},y)$ at each step. Local MAP equals DPS only in Gaussian conditional settings; outside them, local MAP generally does not sample the posterior and can fail to reach the global MAP. We visualize a toy example in Figure 3. Compared to DPS and DAPS, LMAPS is less likely to generate samples in between-mode regions or low-density regions.
4 Local MAP sampling for inverse problem
The primary goal of solving an inverse problem is to recover an unknown image or signal $x_{0}\in\mathbb{R}^{n}$ from a prior distribution, $\pi(x_{0})$, and noisy measurement $y\in\mathbb{R}^{m}$. Mathematically, the unknown signal and the measurements are related by a forward model:
$y=\mathcal{H}(x_{0})+z$ (10)
where $\mathcal{H}(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ (with $m<n$) represents the linear or non-linear forward operator, $z\in\mathbb{R}^{m}$ denotes the noise in the measurement domain. We assume the added noise $z$ is sampled from a Gaussian distribution $\mathcal{N}(0,\sigma_{y}^{2}\mathbb{I})$, where $\sigma_{y}>0$ denotes the noise level. The forward operator and Eq. 10 define the likelihood $p(y\mid x_{0})$ for both the global or local MAP problems in Sec. 3.1.
The final ingredient for constructing a local posterior and solving the resulting MAP problem is the choice of prior $p(x_{0}\mid x_{t})$. While the true transition kernel of a diffusion model prior requires simulation, we can proceed as in previous work (boys2023tweedie, ; song2023pseudoinverse, ) by projecting onto the first two moments using a Gaussian approximation, $p(x_{0}\mid x_{t})\approx\mathcal{N}(x_{0};m_{0\mid t},\Sigma_{0\mid t})$, where $m_{0\mid t}(x_{t}):=\mathbb{E}[x_{0}\mid x_{t}]$. While boys2023tweedie show that $\Sigma_{0\mid t}^{\textsc{tmpd}}(x_{t}):=\mathbb{E}[(x_{0}-m_{0\mid t})(x_{0}-m_{0\mid t})^{T}\mid x_{t}]=\frac{\sigma_{t}^{2}}{\alpha_{t}}\nabla_{x_{t}}m_{0|t}$ to faithfully reflect the covariance of $x_{0}\mid x_{t}$, we will consider flexible choices of $\Sigma_{0|t}$. Finally, the local MAP problem amounts to solving
$x_{0}^{*}=\arg\min(x_{0}-m_{0\mid t})^{T}\Sigma_{0\mid t}^{-1}(x_{0}-m_{0\mid t})+\frac{1}{\sigma_{y}^{2}}\|y-\mathcal{H}(x_{0})\|^{2}.$ (11)
We will develop methodology for approximately solving the local MAP problem for general non-linear inverse problems with differentiable forward operators in Sec. 4.1, before discussing the case of linear inverse problems in Sec. 4.2 or non-differentiable forward operators in Sec. 4.3.
4.1 Approximated solution for nonlinear inverse problems
Isotropic approximation of $\Sigma_{0\mid t}$. For nonlinear $\mathcal{H}(\cdot)$, there is no explicit solution for $x_{0}^{*}$ and it would be more expensive to adopt the moment projection covariance $\Sigma_{0\mid t}=\nabla_{0\mid t}^{\textsc{tmpd}}=\nabla_{x_{t}}m_{0|t}$. We instead choose $\Sigma_{0\mid t}$ based on several observations: (1) the conditional uncertainty cannot be smaller than the injected noise scaled back: $\Sigma_{0\mid t}\preceq\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}\mathbb{I}$; (2) The variance shrinks to 0 as $t$ goes from $T$ to 0; (3) $\Sigma_{0\mid t}$ is is nearly isotropic as $t$ goes to 0. Based on these observations, we assume $\Sigma_{0\mid t}\approx\frac{k}{\text{SNR}}\mathbb{I}$, where $\text{SNR}:=\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}}$. Then $x_{0}^{*}$ in Eq. 11 is reduced to:
$x_{0}^{*}=\arg\min\left\{\frac{\text{SNR}}{k}\|x_{0}-m_{0\mid t}\|^{2}+\frac{1}{\sigma_{y}^{2}}\|y-\mathcal{H}(x_{0})\|^{2}\right\}.$ (12)
Objective Reformulation. In the implementation, the weighting of the two terms in Eq. 12 depends on raw signal-to-noise ratios, which can vary drastically with $t$, which makes it difficult to choose the appropriate learning rate. For analysis and implementation it is convenient to reformulate Eq. 12 in a scale-invariant way. Multiplying the objective by a positive constant (which does not change the minimizer) and introducing parameters $k_{1},k_{2}>0$ such that $2k_{2}/k_{1}^{2}=k/(\alpha_{t}^{2}\sigma_{y}^{2})$ , we obtain the equivalent problem
$x_{0}^{*}=\arg\min\left\{\left(1-\frac{\sigma_{t}^{2}}{\sigma_{t}^{2}+k_{1}^{2}}\right)\frac{1}{2}\|x_{0}-m_{0\mid t}\|^{2}+\frac{\sigma_{t}^{2}}{\sigma_{t}^{2}+k_{1}^{2}}k_{2}\|y-\mathcal{H}(x_{0})\|^{2}\right\}.$ (13)
This reformulation has several advantages:
  • Convex-combination interpretation. The weights can be written as $(1-\mu_{t})$ and $\mu_{t}$ with $\mu_{t}=\sigma_{t}^{2}/(\sigma_{t}^{2}+k_{1}^{2})\in(0,1)$. Thus the cost is a convex combination of the prior and data fidelity terms.
  • Automatic annealing. As $\sigma_{t}^{2}$ decreases over time, $\mu_{t}$ gradually shifts the objective from measurement-driven $\mu_{t}\approx 1$ to prior-driven ($\mu_{t}\approx 0$).
  • Interpretable parameters. The scale $k_{1}$ plays the role of a trust-region parameter balancing prior and measurement, while $k_{2}$ is a scale factor for the consistency loss to the measurement.
  • Numerical stability. Keep weights in $[0,1]$ avoids extreme scaling from SNR values, improving conditioning and optimizer robustness.
In the implementation, we adopt gradient descent to solve $x_{0}^{*}$ in Eq. 13, the algorithm of LMAPS for inverse problems is provided in Algorithm 4.
Relationship to optimization-based methods. Previous optimization-based approaches (song2023solving, ; li2024decoupled, ; zhu2023denoising, ) solve for $x_{0}^{*}$ through the following objective:
$x_{0}^{*}=\arg\min\|x_{0}-m_{0\mid t}\|^{2}+\lambda_{t}\|y-\mathcal{H}(x_{0})\|^{2},$ (14)
where $\lambda_{t}$ is a hyperparameter, often chosen heuristically without a principled basis. These methods can be viewed as special cases of our framework by setting $\Sigma_{0\mid t}=\lambda_{t}\sigma_{y}^{2}\mathbb{I}$ in Eq. 11.
Algorithm 4 Local MAP Sampling (LMAPS) for inverse problems.
1:Input: measurement $y$, forward operator $\mathcal{H}(\cdot)$, pretrained DM $\epsilon_{\theta}(\cdot)$, number of diffusion step $N$, diffusion schedule $\alpha_{t}$ and $\sigma_{t}$, number of gradient updates $K$, objective parameters $k_{1}$, $k_{2}$, learning rate $\eta$.
2:Initialization: $x_{N}\sim\mathcal{N}(0,\mathbb{I})$
3:for $n=N$ to $1$ do
4:  $\hat{x}_{0}\leftarrow[{x_{n}-\sigma_{n}\epsilon_{\theta}(x_{n},n)}]/{\alpha_{n}}$ $\triangleright$ Obtain predicted data
5:  $r\leftarrow\sigma_{n}^{2}/(\sigma_{n}^{2}+k_{1}^{2}+10^{-6})$
6:  Initialization: $x_{0}^{\prime}\leftarrow\hat{x}_{0}$
7:  for $k=K$ to $1$ do
8:   $grad\leftarrow(x_{0}^{\prime}-\hat{x}_{0})(1-r)+rk_{2}\nabla_{x_{0}^{\prime}}\|y-\mathcal{H}(x_{0}^{\prime})\|^{2}$ $\triangleright$ Calculate gradient in Eq. 13
9:   $x_{0}^{\prime}=x_{0}^{\prime}-\eta\cdot grad$
10:  end for
11:  $x_{n-1}\sim\mathcal{N}(\alpha_{n-1}x_{0}^{\prime},\sigma_{n-1}\mathbb{I})$ $\triangleright$ Forward diffusion step
12:end for
13:Output $x_{0}$
While the objectives in Eq. 14 and Eq. 13 are indeed equivalent, we found that empirical performance strongly depends on our objective reformulation and choices of weighting terms as motivated above. Further, our local MAP interpretation provides a probabilistic perspective for these objectives and suggests the connection with TMPD in the case of linear inverse problems, as discussed in Sec. 4.2.
4.2 Exact Solution for Linear Inverse Problems
As discussed in Sec. 3.2, the local MAP solution matches the posterior mean for Gaussian posteriors $p(x_{t}\mid x_{0},y)$ arising from linear inverse problems $p(y\mid x_{0})=\mathcal{N}(Hx_{0},\sigma_{y}^{2}\mathbb{I})$ with a Gaussian assumption on the prior $p(x_{0}\mid x_{t})=\mathcal{N}(x_{0};m_{0\mid t},\Sigma_{0\mid t})$. Solving in closed form for the posterior mean as in (boys2023tweedie, ), we have
$x_{0}^{*}=m_{0\mid t}+\Sigma_{0\mid t}H^{T}(H\Sigma_{0\mid t}H^{T}+\sigma_{y}^{2}\mathbb{I})^{-1}(y-Hm_{0\mid t}).$ (15)
We recover Tweedie Moment-Projected Diffusion (boys2023tweedie, ) as a special case for $\Sigma^{\textsc{tmpd}}_{0\mid t}=\frac{\sigma_{t}^{2}}{\alpha_{t}}\nabla_{x_{t}}m_{0\mid t}(x_{t})$, which is expensive since it requires the gradient with respect to the denoiser $m_{0\mid t}$.
When applying LMAPS to linear inverse problems, we assume $\Sigma_{0|t}=\frac{k}{\text{SNR}_{t}}\mathbb{I}$ as in Sec. 4.1, and optimize with $K$ steps of gradient descent at each timestep despite the availability of the closed form in Eq. 15.
4.3 Addressing inverse problem with non-differentiable operator
Solving the optimization in Eq. 13 requires $\mathcal{H}(\cdot)$ to be differentiable. However, for tasks involving quantization, clipping, or other discrete transforms, computing gradients directly is not feasible. We thus propose an approximation strategy to calculate the gradient term $\nabla_{x_{0}}\|y-\mathcal{H}(x_{0})\|^{2}$.
First, if $\mathcal{H}(\cdot)$ is differentiable, the gradient of the residual term with respect to the measurement $\|y-\mathcal{H}(x_{0})\|^{2}$ is given by:
$\nabla_{x_{0}}\|y-\mathcal{H}(x_{0})\|^{2}\approx 2J_{\mathcal{H}}(x_{0})^{T}\big(\mathcal{H}(x_{0})-y\big),$ (16)
where $J_{\mathcal{H}}(x_{0})$ denotes the Jacobian of $\mathcal{H}$ at $x_{0}$. For non-differentiable $\mathcal{H}(\cdot)$, we suppose there exists a surrogate operator $\mathcal{H}^{\prime}(\cdot)$ that is differentiable and sufficiently close to $\mathcal{H}(\cdot)$ in the sense that the error $\mathcal{H}(x_{0})-\mathcal{H}^{\prime}(x_{0})$ can be neglected. In this case, the gradient term can be approximated by:
$\nabla_{x_{0}}\|y-\mathcal{H}(x_{0})\|^{2}\approx 2J_{\mathcal{H}^{\prime}}(x_{0})^{T}\big(\mathcal{H}(x_{0})-y\big).$ (17)
JPEG restoration and quantization. A representative example arises in JPEG artifact removal, where $\mathcal{H}$ corresponds to quantization, which is inherently non-differentiable. In this case, we can take a simple surrogate $\mathcal{H}^{\prime}(x_{0})=x_{0}$, so that $J_{\mathcal{H}^{\prime}}(x_{0})=\mathbb{I}$. The approximation then reduces to:
$\nabla_{x_{0}}\|y-\mathcal{H}(x_{0})\|^{2}\approx 2\big(\mathcal{H}(x_{0})-y\big).$ (18)
This surrogate-gradient approach provides a novel way to solve inverse problems with non-differentiable operators using the diffusion framework.
5 Experiments
5.1 Experimental setup
Inverse problems. We evaluate our method on image restoration and scientific inverse problems. For linear image restoration, we consider (1) super-resolution, (2) Gaussian deblurring, (3) motion deblurring, (4) inpainting (with a box mask), and (5) inpainting (with a 70% random mask). For nonlinear image restoration, we consider (1) phase retrieval, (2) high dynamic range (HDR) reconstruction, (3) nonlinear deblurring, (4) JPEG restoration, (5) quantization, where HDR, JPEG restoration and quantization are nonlinear inverse problems with non-differentiable operators. For scientific inverse problems, we adopt the benchmark from InverseBench (zheng2025inversebench, ), which includes Linear Inverse Scattering (LIS), Compressed sensing MRI (CS-MRI) and Black Hole Imaging. More details are provided in the App. C.
Dataset. For image restoration, we evaluated our method on FFHQ (karras2019style, ) $256\times 256$ and ImageNet $256\times 256$ datasets (deng2009imagenet, ). Following DAPS, we test the same subset of 100 images for both datasets. For scientific inverse problems, we adopt the same dataset as InverseBench (zheng2025inversebench, ).
Pretrained models. For image restoration tasks, we utilize the pre-trained checkpoint (chung2022diffusion, ) on the FFHQ dataset and the pre-trained checkpoint (dhariwal2021diffusion, ) on the ImageNet dataset. For scientific inverse problems, we adopt the pre-trained checkpoints from InverseBench.
[…]
Metrics. For image restoration tasks, we report Peak Signal-to-Noise Ratio (PSNR), Structural SIMilarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) (zhang2018unreasonable, ). For scientific inverse problems, we primarily present PSNR in the main text, while additional task-specific metrics are provided in the App. D.
Figure 4: Ablation study on optimization steps vs. diffusion steps (NFEs) for Gaussian Deblurring.
Table 1: Quantitative evaluation of solving image restoration FFHQ (left) and ImageNet (right), with Gaussian noise (known variance, $\sigma_{y}=0.05$): 5 linear and 5 nonlinear tasks (3 non-differentiable). Results are reported as mean PSNR, SSIM, and LPIPS across 100 validation images; for nonlinear tasks, both mean and standard deviation are provided due to higher instability. Best results are highlighted in bold. LMAPS achieves superior performance on most tasks, with computational cost less than or equal to DAPS. For phase retrieval, DAPS and LMAPS select the best result from 4 runs for each image.
Task Method FFHQ ImageNet
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
SR $4\times$ DPS $25.86$ $0.753$ $0.269$ $21.13$ $0.489$ $0.361$
DDRM $26.58$ $0.782$ $0.282$ $22.62$ $0.521$ $0.324$
DDNM $28.03$ $0.795$ $0.197$ $23.96$ $0.604$ $0.475$
DCDP $28.66$ $0.807$ $0.178$
FPS-SMC $28.42$ $0.813$ $0.204$ $24.82$ $0.703$ $0.313$
DiffPIR $26.64$ $0.260$ $23.18$ $0.371$
DAPS $29.07$ $0.818$ $0.177$ $25.89$ $0.694$ $0.276$
SITCOM $30.55$ $0.864$ $\mathbf{0.154}$ $\mathbf{27.07}$ $\mathbf{0.746}$ $\mathbf{0.228}$
LMAPS $\mathbf{30.74}$ $\mathbf{0.869}$ $0.165$ $26.72$ $0.739$ $0.242$
Inpaint (Box) DPS $22.51$ $0.792$ $0.209$ $18.94$ $0.722$ $0.257$
DDRM $22.26$ $0.801$ $0.207$ $18.63$ $0.733$ $0.254$
DDNM $24.47$ $0.837$ $0.235$ $21.64$ $0.748$ $0.319$
DCDP $23.89$ $0.760$ $0.163$
FPS-SMC $24.86$ $0.823$ $0.146$ $\mathbf{22.16}$ $0.726$ $0.208$
DAPS $24.07$ $0.814$ $0.133$ $21.43$ $0.725$ $0.214$
SITCOM $24.95$ $0.849$ $0.131$ $19.72$ $0.784$ $\mathbf{0.164}$
LMAPS $\mathbf{25.17}$ $\mathbf{0.876}$ $\mathbf{0.108}$ $21.25$ $\mathbf{0.803}$ $0.204$
Inpaint (Random) DPS $25.46$ $0.823$ $0.203$ $23.52$ $0.745$ $0.297$
DDNM $29.91$ $0.817$ $0.121$ $\mathbf{31.16}$ $0.841$ $0.191$
DCDP $30.69$ $0.842$ $0.142$
FPS-SMC $28.21$ $0.823$ $0.261$ $24.52$ $0.701$ $0.316$
DAPS $31.12$ $0.844$ $0.098$ $28.44$ $0.775$ $0.135$
SITCOM $33.96$ $0.928$ $0.082$ $29.74$ $0.855$ $0.115$
LMAPS $\mathbf{34.51}$ $\mathbf{0.938}$ $\mathbf{0.066}$ $30.59$ $\mathbf{0.876}$ $\mathbf{0.100}$
Gaussian Deblurring DPS $25.87$ $0.764$ $0.219$ $20.31$ $0.598$ $0.397$
DDRM $24.93$ $0.732$ $0.239$ $21.26$ $0.564$ $0.443$
DCDP $27.50$ $0.699$ $0.304$
FPS-SMC $26.54$ $0.773$ $0.253$ $23.91$ $0.601$ $0.387$
DiffPIR $27.36$ $0.236$ $22.80$ $0.355$
DAPS $29.19$ $0.817$ $0.165$ $26.15$ $0.684$ $0.253$
SITCOM $29.93$ $0.846$ $0.172$ $26.39$ $0.716$ $0.260$
LMAPS $\mathbf{30.88}$ $\mathbf{0.867}$ $\mathbf{0.158}$ $\mathbf{26.65}$ $\mathbf{0.727}$ $\mathbf{0.250}$
Motion Deblurring DPS $24.52$ $0.801$ $0.246$ $18.96$ $0.629$ $0.423$
DCDP $25.08$ $0.512$ $0.364$
FPS-SMC $27.39$ $0.826$ $0.227$ $24.52$ $0.647$ $0.326$
DiffPIR $26.57$ $0.255$ $24.01$ $0.366$
DAPS $29.66$ $0.847$ $0.157$ $27.86$ $0.766$ $0.196$
SITCOM $29.36$ $0.840$ $0.185$ $26.76$ $0.746$ $0.242$
LMAPS $\mathbf{32.62}$ $\mathbf{0.902}$ $\mathbf{0.117}$ $\mathbf{28.42}$ $\mathbf{0.796}$ $\mathbf{0.204}$
Phase Retrieval DPS $17.64_{\pm 2.97}$ $0.441_{\pm 0.129}$ $0.410_{\pm 0.090}$ $16.81_{\pm 3.61}$ $0.427_{\pm 0.143}$ $0.447_{\pm 0.099}$
RED-diff $15.60_{\pm 4.48}$ $0.398_{\pm 0.195}$ $0.596_{\pm 0.092}$ $14.98_{\pm 3.75}$ $0.386_{\pm 0.057}$ $0.536_{\pm 0.129}$
DAPS $30.63_{\pm 3.13}$ $0.851_{\pm 0.072}$ $0.139_{\pm 0.060}$ $21.39_{\pm 6.59}$ $0.473_{\pm 0.226}$ $0.372_{\pm 0.166}$
LMAPS $\mathbf{31.56_{\pm 3.02}}$ $\mathbf{0.867_{\pm 0.057}}$ $\mathbf{0.126_{\pm 0.052}}$ $\mathbf{22.86_{\pm 7.50}}$ $\mathbf{0.596_{\pm 0.267}}$ $\mathbf{0.313_{\pm 0.176}}$
Nonlinear Deblurring DPS $23.39_{\pm 2.01}$ $0.263_{\pm 0.082}$ $0.278_{\pm 0.060}$ $22.49_{\pm 3.20}$ $0.591_{\pm 0.101}$ $0.306_{\pm 0.081}$
RED-diff $\mathbf{30.86_{\pm 0.51}}$ $0.795_{\pm 0.028}$ $0.160_{\pm 0.034}$ $\mathbf{30.07_{\pm 1.41}}$ $0.754_{\pm 0.023}$ $0.211_{\pm 0.083}$
DCDP $27.92_{\pm 2.64}$ $0.779_{\pm 0.067}$ $0.183_{\pm 0.051}$
DAPS $28.29_{\pm 1.77}$ $0.783_{\pm 0.036}$ $0.155_{\pm 0.032}$ $27.73_{\pm 3.23}$ $0.724_{\pm 0.048}$ $0.169_{\pm 0.056}$
SITCOM $29.19_{\pm 2.35}$ $0.785_{\pm 0.093}$ $0.190_{\pm 0.014}$ $28.55_{\pm 3.87}$ $\mathbf{0.798_{\pm 0.092}}$ $\mathbf{0.149_{\pm 0.050}}$
LMAPS $29.93_{\pm 1.83}$ $\mathbf{0.855_{\pm 0.035}}$ $\mathbf{0.150_{\pm 0.034}}$ $28.03_{\pm 3.62}$ $0.774_{\pm 0.099}$ $0.183_{\pm 0.065}$
High Dynamic Range DPS $22.73_{\pm 6.07}$ $0.591_{\pm 0.141}$ $0.264_{\pm 0.156}$ $19.23_{\pm 2.52}$ $0.582_{\pm 0.082}$ $0.503_{\pm 0.106}$
DAPS $27.12_{\pm 3.53}$ $0.752_{\pm 0.041}$ $0.162_{\pm 0.072}$ $26.30_{\pm 4.10}$ $0.717_{\pm 0.067}$ $0.175_{\pm 0.107}$
SITCOM $28.02_{\pm 3.28}$ $0.812_{\pm 0.108}$ $0.174_{\pm 0.081}$ $25.59_{\pm 3.66}$ $0.170_{\pm 0.141}$ $0.198_{\pm 0.177}$
LMAPS $\mathbf{28.87_{\pm 3.39}}$ $\mathbf{0.884_{\pm 0.082}}$ $\mathbf{0.141_{\pm 0.074}}$ $\mathbf{27.02_{\pm 4.00}}$ $\mathbf{0.860_{\pm 0.096}}$ $\mathbf{0.158_{\pm 0.090}}$
[…]
LMAPS $\mathbf{27.25_{\pm 1.37}}$ $\mathbf{0.814_{\pm 0.045}}$ $\mathbf{0.260_{\pm 0.043}}$ $\mathbf{24.96_{\pm 2.46}}$ $\mathbf{0.703_{\pm 0.124}}$ $\mathbf{0.340_{\pm 0.089}}$
[…]
LMAPS $\mathbf{29.51_{\pm 1.14}}$ $\mathbf{0.844_{\pm 0.467}}$ $\mathbf{0.229_{\pm 0.474}}$ $\mathbf{26.92_{\pm 2.25}}$ $\mathbf{0.748_{\pm 0.114}}$ $\mathbf{0.307_{\pm 0.099}}$
5.2 Main results
Ablation studies. Figure 4 presents ablation studies on optimization steps across different diffusion steps. The best performance is typically observed at NFE = 200–500, where increasing the number of optimization steps per diffusion step. yields notable improvements. Compared to the baseline SITCOM (600 NFEs with gradient computation through the U-Net), LMAPS attains similar performance while requiring substantially fewer computational resources. We report runtime comparisons for various methods on Deblurring task in Table 3 (App. B).
Image restoration. In Table 1, we present quantitative results for image restoration tasks on FFHQ and ImageNet datasets. The table covers 10 tasks, 3 restoration quality metrics, and 2 datasets, totaling 60 results. LMAPS achieves the best performance in 49 out of 60 cases. Generally, LMAPS demonstrates superior performance than DAPS for most of the tasks with less computational cost. LMAPS improves >> 2dB PSNR across motion deblurring, JPEG restoration and quantization tasks.
Scientific inverse problems. In Table 2, we report quantitative results of solving scientific inverse problems: Linear Inverse Scattering (LIS), CS-MRI, Black Hole Imaging. LMAPS demonstrates the best PSNR across all tasks, improved more than 1.5 dB PSNR for 3 LIS tasks.
Table 2: Quantitative evaluation of solving scientific inverse problems is conducted using PSNR as the evaluation metric. The tasks include: (i) three LIS settings with different numbers of receivers (NR = 360, 180, 60); (ii) four CS-MRI settings with varying subsampling ratios ($4\times$, $8\times$) and measurement types (noiseless and raw); and (iii) three Black Hole Imaging settings with different observation time ratios ($3\%$, $\%, 1001$$0\%$).
Method LIS CS-MRI Black Hole
NR=$360$ NR=$180$ NR=$60$ $4\times$ noiseless $4\times$ raw $8\times$ noiseless $8\times$ raw $0\% 1010\%$ $3\%$
DDRM $32.13$ $28.08$ $20.44$
DDNM $26.28$ $35.02$ $29.24$
[…]
DPS $32.06$ $31.80$ $27.37$ $26.13$ $25.83$ $20.82$ $23.00$ $25.86$ $24.36$ $24.20$
LGD $27.90$ $27.84$ $20.49$ $21.22$ $22.08$ $22.51$
DiffPIR $34.24$ $34.01$ $26.32$ $28.31$ $27.60$ $26.78$ $26.26$ $25.01$ $23.84$ $24.12$
PnP-DM $33.94$ $31.82$ $24.72$ $31.80$ $27.62$ $29.33$ $25.28$ $26.07$ $24.57$ $24.25$
DAPS $34.64$ $33.16$ $25.88$ $31.48$ $28.61$ $29.01$ $27.10$ $25.60$ $23.99$ $23.54$
RED-diff $36.56$ $35.41$ $27.07$ $29.36$ $28.71$ $26.76$ $27.33$ $23.77$ $22.53$ $20.74$
FPS $33.24$ $29.62$ $21.32$
MCG-diff $30.94$ $28.06$ $21.00$
LMAPS $\mathbf{38.07}$ $\mathbf{37.19}$ $\mathbf{30.75}$ $\mathbf{32.83}$ $\mathbf{28.77}$ $\mathbf{30.50}$ $\mathbf{27.43}$ $\mathbf{26.79}$ $\mathbf{24.83}$ $\mathbf{24.66}$
6 Related work
Recent advances in conditional generation have led to breakthroughs in text-to-image synthesis, semantic editing, and domain-specific applications such as image-to-image translation and controlled signal reconstruction (song2023loss, ; ye2024tfg, ; skreta2025feynman, ; singhal2025general, ; zheng2023guided, ). These methods have been especially impactful in solving inverse problems, including image restoration and scientific reconstruction tasks (wang2022zero, ; zheng2025inversebench, ). A wide range of approaches have been developed, spanning linear projection methods (wang2022zero, ; kawar2022denoising, ; zhang2025measurement, ; dou2024diffusion, ), Monte Carlo sampling (wu2023practical, ; phillips2024particle, ), variational inference (feng2023score, ; mardani2023variational, ; janati2024divide, ), and optimization-based strategies (song2023solving, ; zhu2023denoising, ; li2024decoupled, ; wang2024dmplug, ; alkhouri2024sitcom, ; he2023manifold, ).
Among these, Diffusion Posterior Sampling (DPS) and its variants (zhang2025improving, ; chung2022diffusion, ; song2023loss, ; yu2023freedom, ; rout2024beyond, ; yang2024guidance, ; bansal2023universal, ; boys2023tweedie, ; song2023pseudoinverse, ; ho2022classifier, ) have gained wide adoption due to their strong empirical performance and interpretability, as they directly sample from the posterior distribution $p(x_{0}\mid y)$. More recently, attention has shifted toward maximum a posteriori (MAP) estimation with diffusion priors. xu2025rethinking argued that DPS is in fact more consistent with the principle of MAP estimation rather than true posterior sampling, although their proposed sampling algorithm differs from ours. Finally, gutha2025inverse proposed sampling from the global MAP solution, $\arg\max p(x_{0}\mid y)$, though their approach is largely restricted to linear inverse problems.
7 Conclusion
We presented Local MAP Sampling (LMAPS), a new inference framework that that iteratively solves local maximum-a-posteriori subproblems along the diffusion trajectory. By introducing a principled covariance approximation, an objective reformulation, and a gradient strategy for non-differentiable operators, LMAPS provides both theoretical clarity and practical effectiveness. Experiments across diverse image restoration and scientific inverse problems show that LMAPS consistently improves reconstruction quality, particularly on challenging tasks such as Box Inpainting, Phase Retrieval, JPEG restoration, and HDR.
Future work. In Bayesian inference, the global MAP plays a critical role and offers valuable insights when contrasted with posterior sampling. Yet its utility has been largely overlooked, and efficiently solving the global MAP with diffusion priors remains an open challenge. While MAP may reduce output diversity by concentrating on a single mode, it provides higher certainty and potentially better alignment with observed data. Advancing in this direction could enable more probable reconstructions and make significant contributions to solving inverse problems.
Appendix A Gaussian mixture toy example
To gain intuition about posterior mean and MAP estimates in diffusion models, we consider a tractable toy prior $\pi_{0}(x_{0})$ given by a Gaussian mixture:
$\pi_{0}(x_{0})=\sum_{k=1}^{K}\pi_{k}\,\mathcal{N}(x_{0};\,\mu_{k},\Sigma_{k}),$ (19)
where $\pi_{k}>0$ and $\sum_{k}\pi_{k}=1$.
Forward kernel.
As in the unconditional diffusion model, the forward corruption is
$p_{t}(x_{t}\mid x_{0})=\mathcal{N}(x_{t};\,\alpha_{t}x_{0},\sigma_{t}^{2}I).$ (20)
Thus the marginal $p_{t}(x_{t})=\int p_{t}(x_{t}\mid x_{0})\,\pi_{0}(x_{0})\,dx_{0}$ is itself a Gaussian mixture.
Posterior distribution.
By Bayes’ rule,
$p(x_{0}\mid x_{t})\propto p_{t}(x_{t}\mid x_{0})\,\pi_{0}(x_{0}).$ (21)
Conditioned on mixture component $k$, the posterior remains Gaussian:
$p(x_{0}\mid x_{t},k)$ $=\mathcal{N}(x_{0};\,m_{k},S_{k}),$ (22)
$S_{k}$ $=\Big(\Sigma_{k}^{-1}+\tfrac{\alpha_{t}^{2}}{\sigma_{t}^{2}}I\Big)^{-1},$ (23)
$m_{k}$ $=S_{k}\!\left(\Sigma_{k}^{-1}\mu_{k}+\tfrac{\alpha_{t}}{\sigma_{t}^{2}}x_{t}\right).$ (24)
The responsibilities are
$r_{k}(x_{t})=\frac{\pi_{k}\,\mathcal{N}(x_{t};\,\alpha_{t}\mu_{k},\,\alpha_{t}^{2}\Sigma_{k}+\sigma_{t}^{2}I)}{\sum_{j}\pi_{j}\,\mathcal{N}(x_{t};\,\alpha_{t}\mu_{j},\,\alpha_{t}^{2}\Sigma_{j}+\sigma_{t}^{2}I)}.$ (25)
Hence the full posterior is itself a Gaussian mixture:
$p(x_{0}\mid x_{t})=\sum_{k=1}^{K}r_{k}(x_{t})\,\mathcal{N}(x_{0};\,m_{k},S_{k}).$ (26)
Posterior mean.
The ideal denoiser in this case has a closed form:
$m_{0\mid t}(x_{t}):=\mathbb{E}[x_{0}\mid x_{t}]=\sum_{k=1}^{K}r_{k}(x_{t})\,m_{k}.$ (27)
This estimator is mode-averaging, and may fall between mixture modes.
Local MAP.
Each component posterior has its mode at $m_{k}$. A local MAP predictor is obtained by selecting the component with the highest posterior peak density,
$k^{\star}(x_{t})=\arg\max_{k}\;\frac{r_{k}(x_{t})}{\sqrt{(2\pi)^{d}\det S_{k}}},\qquad x_{0}^{*}(t,x_{t})=m_{k^{\star}(x_{t})}.$ (28)
Unlike the posterior mean, this estimate is mode-seeking and stays in high-density regions.
DDIM iteration.
Replacing the generic denoiser $m_{0\mid t}(x_{t})$ in the DDIM update with either the posterior mean, local MAP, or a posterior sample yields three distinct variants of the reverse process:
$x_{t-\Delta t}$ $=g(m_{0\mid t}(x_{t}),\,x_{t},\epsilon)$ (posterior mean) (29)
$x_{t-\Delta t}$ $=g(x_{0}^{*}(t,x_{t}),\,x_{t},\epsilon)$ (local MAP) (30)
$x_{t-\Delta t}$ $=g(\tilde{x}_{0},\,x_{t},\epsilon),\;\tilde{x}_{0}\sim p(x_{0}\mid x_{t})$ $\text{(local sampling)}.$ (31)
This toy setup makes explicit the distinction between mean-based denoising and MAP-based denoising.
Appendix B Sampling Efficiency
We present a comparison of sampling times among LMAPS, DAPS, and SITCOM. Among them, SITCOM and DAPS achieve the third- and second-best results, respectively, while LMAPS demonstrates the best performance with lower computation time.
Table 3: Sampling time of LMAPS on Deblurring tasks with FFHQ 256. The non-parallel single-image sampling time on the FFHQ 256 dataset using one NVIDIA A6000 GPU. NFE refers to diffusion timesteps, while optimization steps refer to inner loop optimizations in respective methods.
Configuration ODE Steps Optimization Steps NFE Second/Image PSNR
DAPS 5 100 200 110 29.19
SITCOM 30 600 73 29.93
DPS 1000 138 25.87
MAPS 100 200 61 30.88
10 100 6 30.00
100 20 6 30.64
20 100 6 30.43
20 20 2 29.91
Appendix C Experiment details
C.1 Dataset details
For scientific inverse problems, we adopt fluorescence microscopy images from InverseBench [40] on linear inverse scattering tasks, General Relativistic MagnetoHydroDynamic (GRMHD) [20] on black hole imaging, and multi-coil raw $k$-space data from the fastMRI knee dataset [36] on CS-MRI.
C.2 Inverse problem details
Baselines from DAPS [37]. For image restoration tasks include: (1) super-resolution, (2) Gaussian deblurring, (3) motion deblurring, (4) inpainting (with a box mask), and (5) inpainting (with a 70% random mask), (6) phase retrieval, (7) high dynamic range (HDR) reconstruction, (8) nonlinear deblurring, we follow the same experimental setup as in DAPS.
InverseBench [40]. For scientific inverse problems, we adopt the setting introduced in InverseBench.
JPEG Restoration. We address JPEG restoration with quality factors of $\mathrm{QF}=5$.
Quantization. We model quantization by discretizing the measurement into $2^{n_{\text{bits}}}$ uniformly spaced levels. Formally, the forward operator is defined as
$\mathcal{H}(x)=\frac{\left\lfloor x\cdot(2^{n_{\text{bits}}}-1)+0.5\right\rfloor}{2^{n_{\text{bits}}}-1},$ (32)
which rounds the input $x$ to the nearest quantization level. In this work, we focus on the challenging case of 2-bit quantization, where only four distinct measurement levels are available, significantly reducing precision and making accurate reconstruction more difficult.
C.3 Baseline details
For SITCOM [1], we use the hyperparameter configuration recommended in the original paper, with $N=20$ and $K=30$, resulting in $600$ NFEs and requiring gradient computation with respect to the U-Net.
[…]
Other baselines we adopt the same reported results as in DAPS [37] and InverseBench [40].
C.4 Complete List of hyper-parameters
We provide complete lost of hyper-paramers of LMAPS for different inverse problems in Table 4.
Tasks Annealing step Gradient step Learning rate $\eta$ $k_{1}$ $k_{2}$
SR $4\times$ 200 100 0.05 0.15 20
Inpaint (Box) 200 100 0.02 0.5 50
Inpaint (Random) 200 100 0..01 0.22 100
Gaussian Deblurring 200 100 0.01 0.22 100
Motion Deblurring 200 100 0.01 0.25 100
Phase Retrieval 200 100 0.1 10 0.3
Nonlinear Deblurring 200 100 0.02 0.05 1
High Dynamic Range 200 100 0.04 0.2 10
JPEG Restoration 200 100 0.2 0.5 5
Quantization 200 20 0.2 0.5 5
LIS (NR=360) 200 50 1 0 5000
LIS (NR=180) 200 50 1 0 10000
LIS (NR=60) 200 50 1 0 30000
CS-MRI ($4\times$, noiseless) 200 100 0.01 0 100
CS-MRI ($4\times$, raw) 200 100 0.01 0.4 150
CS-MRI ($8\times$, noiseless) 200 100 0.01 0.4 150
CS-MRI ($8\times$, raw) 200 100 0.01 0.4 150
[…]
Table 4: Complete List of hyper-parameters of LMAPS for different inverse problems.
Appendix D Additional experiment results
D.1 Scientific inverse problems
We present additional evaluation metrics on linear inverse scattering in Table 5, compressed sensing MRI in Table 6, and black hole imaging in Table 7.
Table 5: Results on Linear inverse scattering. PSNR and SSIM of different algorithms on linear inverse scattering. Noise level $\sigma_{y}=10^{-4}$.
Number of receivers 360 180 60
Methods PSNR SSIM PSNR SSIM PSNR SSIM
Traditional
FISTA-TV 32.126 (2.139) 0.979 (0.009) 26.523 (2.678) 0.914 (0.040) 20.938 (2.513) 0.709 (0.103)
PnP diffusion prior
DDRM 32.598 (1.825) 0.929 (0.012) 28.080 (1.516) 0.890 (0.019) 20.436 (1.210) 0.545 (0.037)
DDNM 36.381 (1.098) 0.935 (0.017) 35.024 (0.993) 0.895 (0.027) 29.235 (3.376) 0.917 (0.022)
IIGDM 27.925 (3.211) 0.889 (0.072) 26.412 (3.430) 0.816 (0.114) 20.074 (2.608) 0.540 (0.198)
DPS 32.061 (2.163) 0.846 (0.127) 31.798 (2.163) 0.862 (0.123) 27.372 (3.415) 0.813 (0.133)
LGD 27.901 (2.346) 0.812 (0.037) 27.837 (3.031) 0.803 (0.034) 20.491 (3.031) 0.552 (0.077)
DiffPIR 34.241 (2.310) 0.988 (0.006) 34.010 (2.269) 0.987 (0.006) 26.321 (3.272) 0.918 (0.028)
PnP-DM 33.914 (2.054) 0.988 (0.006) 31.817 (2.073) 0.981 (0.008) 24.715 (2.874) 0.909 (0.046)
DAPS 34.641 (1.693) 0.957 (0.006) 33.160 (1.704) 0.944 (0.009) 25.875 (3.110) 0.885 (0.030)
RED-diff 36.556 (2.292) 0.981 (0.005) 35.411 (2.166) 0.984 (0.004) 27.072 (3.330) 0.935 (0.037)
FPS 33.242 (1.602) 0.870 (0.026) 29.624 (1.651) 0.710 (0.040) 21.323 (1.445) 0.460 (0.030)
MCG-diff 30.937 (1.964) 0.751 (0.029) 28.057 (1.672) 0.631 (0.042) 21.004 (1.571) 0.445 (0.028)
LMAPS 38.074 (1.905) 0.994 (0.001) 37.188 (1.815) 0.990 (0.001) 30.759 (3.539) 0.967 (0.211)
Table 6: Results on compressed sensing MRI. Mean and standard deviation are reported over 94 test cases. Underline: the best across all methods. Bold: the best across PnP DM methods.
Methods $\times 4$ Simulated (noiseless) $\times 4$ Raw $\times 8$ Simulated (noiseless) $\times 8$ Raw
PSNR $\uparrow$ SSIM $\uparrow$ Data misfit $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ Data misfit $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ Data misfit $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ Data misfit $\downarrow$
Traditional
[…]
TV 27.03 (1.635) 0.518 (0.123) 5.748 (1.283) 26.22 (1.578) 0.509 (0.123) 32.269 (15.414) 24.12 (1.900) 0.432 (1.112) 5.087 (1.049) 23.70 (1.857) 0.427 (0.112) 23.048 (10.854)
End-to-end
Residual UNet 32.27 (1.810) 0.808 (0.080) 31.70 (1.970) 0.785 (0.095) 29.75 (1.675) 0.750 (0.088) 29.36 (1.746) 0.733 (0.100)
E2E-VarNet 33.40 (2.097) 0.836 (0.079) 31.71 (2.540) 0.756 (0.102) 30.67 (1.761) 0.769 (0.085) 30.45 (1.940) 0.736 (0.103)
PnP diffusion prior
CSGM 28.78 (6.173) 0.710 (0.147) 1.518 (0.433) 25.17 (6.246) 0.582 (0.167) 31.642 (15.382) 26.15 (6.383) 0.625 (0.158) 1.142 (1.078) 21.17 (8.314) 0.425 (0.192) 22.088 (10.740)
ScoreMRI 25.97 (1.681) 0.468 (0.087) 10.828 (1.731) 25.60 (1.618) 0.463 (0.086) 33.697 (15.209) 25.20 (1.526) 0.405 (0.079) 8.360 (1.381) 24.74 (1.481) 0.403 (0.080) 24.028 (10.663)
RED-diff 29.36 (7.710) 0.733 (0.131) 0.509 (0.077) 28.71 (2.755) 0.626 (0.126) 31.591 (15.368) 26.76 (6.969) 0.647 (0.124) 0.485 (0.068) 27.33 (2.441) 0.563 (0.117) 22.336 (10.838)
DiffPIR 28.31 (1.598) 0.632 (0.107) 10.545 (2.466) 27.60 (1.470) 0.624 (0.111) 34.015 (15.522) 26.78 (1.556) 0.588 (0.113) 7.787 (1.741) 26.26 (1.458) 0.590 (0.113) 24.208 (10.922)
DPS 26.13 (4.247) 0.620 (0.105) 9.092 (2.925) 25.83 (2.197) 0.548 (0.116) 35.009 (15.967) 22.82 (4.777) 0.536 (0.111) 6.737 (1.928) 23.00 (3.205) 0.507 (0.109) 24.842 (11.263)
DAPS 31.48 (1.988) 0.762 (0.089) 1.566 (0.390) 28.61 (2.197) 0.689 (0.102) 31.115 (15.497) 29.01 (1.712) 0.681 (0.098) 1.280 (0.301) 27.10 (2.034) 0.629 (0.107) 22.729 (10.926)
PnP-DM 31.80 (3.473) 0.780 (0.096) 4.701 (0.675) 27.62 (3.425) 0.679 (0.117) 32.261 (15.169) 29.33 (3.081) 0.704 (0.105) 3.421 (0.504) 25.28 (3.102) 0.607 (0.117) 22.879 (10.712)
LMAPS 32.83 (2.581) 0.740 (0.117) 3.500 (0.544) 28.77 (1.813) 0.656 (0.102) 32.476 (15.303) 30.50 (2.181) 0.660 (0.116) 2.565 (0.399) 27.43 (1.689) 0.600 (0.109) 23.021 (10.804)
Table 7: Results on black hole imaging. PSNR and Chi-squared of different algorithms on black hole imaging. Gain and phase noise and thermal noise are added based on EHT library.
Methods 3% 10% 100%
PSNR Blur PSNR $\tilde{\chi}^{2}_{cp}$ $\tilde{\chi}^{2}_{logca}$ PSNR Blur PSNR $\tilde{\chi}^{2}_{cp}$ $\tilde{\chi}^{2}_{logca}$ PSNR Blur PSNR $\tilde{\chi}^{2}_{cp}$ $\tilde{\chi}^{2}_{logca}$
Traditional
SMILI 18.51 (1.39) 23.08 (2.12) 1.478 (0.428) 4.348 (3.827) 20.85 (2.90) 25.24 (3.86) 1.209 (0.169) 21.788 (12.491) 22.67 (3.13) 27.79 (4.02) 1.878 (0.952) 17.612 (10.299)
EHT-Imaging 21.72 (3.39) 25.66 (5.04) 1.507 (0.485) 1.695 (0.539) 22.67 (3.46) 26.66 (3.93) 1.166 (0.156) 1.240 (0.205) 24.28 (3.63) 28.57 (4.52) 1.251 (0.250) 1.259 (0.316)
PnP diffusion prior
DPS 24.20 (3.72) 30.83 (5.58) 8.024 (24.336) 5.007 (5.750) 24.36 (3.72) 30.79 (5.75) 13.052 (43.087) 6.614 (26.789) 25.86 (3.90) 32.94 (6.19) 8.759 (37.784) 5.456 (24.185)
LGD 22.51 (3.76) 28.50 (5.49) 15.825 (16.838) 12.862 (12.663) 22.08 (3.75) 27.48 (5.09) 10.775 (21.684) 13.375 (56.397) 21.22 (3.64) 26.06 (4.98) 13.239 (17.231) 13.233 (39.107)
RED-diff 20.74 (2.62) 26.10 (3.35) 6.713 (6.925) 9.128 (19.052) 22.53 (3.02) 27.67 (4.53) 2.488 (2.925) 4.916 (13.221) 23.77 (4.13) 29.13 (6.22) 1.853 (0.938) 2.050 (2.361)
PnPDM 24.25 (3.45) 30.49 (4.93) 2.201 (1.352) 1.668 (0.551) 24.57 (3.47) 30.80 (5.22) 1.433 (0.417) 1.336 (0.478) 26.07 (3.70) 32.88 (6.02) 1.311 (0.195) 1.199 (0.221)
DAPS 23.54 (3.28) 29.48 (4.88) 3.647 (3.287) 2.329 (1.354) 23.99 (3.56) 30.16 (5.13) 1.545 (0.705) 2.253 (9.903) 25.60 (3.64) 32.78 (5.68) 1.300 (0.324) 1.229 (0.532)
DiffPIR 24.12 (3.25) 30.45 (4.88) 14.085 (14.105) 10.545 (8.860) 23.84 (3.59) 30.04 (5.03) 5.374 (3.733) 5.205 (5.556) 25.01 (4.64) 31.86 (6.56) 3.271 (1.623) 2.970 (1.202)
LMAPS 24.66 (4.02) 29.94 (5.17) 1.497 (0.394) 4.695 (1.420) 24.84 (3.695) 29.98 (5.144) 1.671 (0.521) 4.460 (1.555) 26.79 (3.78) 32.95 (5.41) 1.512 (0.474) 4.622 (1.455)
D.2 Additional visualization
Additional visualization are presented in Figs. 5, 6, 7, 8, 9, 10, 11, 12, 13, 14.
Appendix E Licenses
FFHQ Dataset. We use the Flickr-Faces-HQ (FFHQ) dataset released by NVIDIA under the Creative Commons BY-NC-SA 4.0 license. The dataset is intended for non-commercial research purposes only. More details are available at: https://github.com/NVlabs/ffhq-dataset.
ImageNet Dataset. The ImageNet dataset is used under the terms of its academic research license. Access requires agreement to ImageNet’s data use policy, and redistribution is not permitted. More information is available at: https://image-net.org/download.
Figure 5: Visualization of CS-MRI restoration ($4\times$ raw).
Figure 6: Visualization of Linear Inverse Scattering (Number of receivers = 60).
Figure 7: Visualization for solving JPEG restoration (QF=5, $\sigma_{y}=0.05$). Top: degraded images; bottom: generated images.
Figure 8: Visualization for solving Quantization (2 bit). Top: degraded images; bottom: generated images.
Figure 9: Visualization for solving Inpaint (Box). Top: ground truth; middle: degraded images; bottom: generated images.
Figure 10: Visualization for solving HDR. Top: ground truth; middle: degraded images; bottom: generated images.
Figure 11: Visualization for solving Deblurring. Top: ground truth; middle: degraded images; bottom: generated images.
Figure 12: Visualization for solving Super-Resolution. Top: ground truth; middle: degraded images; bottom: generated images.
Figure 13: Visualization for solving Nonlinear Deblurring. Top: ground truth; middle: degraded images; bottom: generated images.
Figure 14: Visualization for solving Phase retrieval.

### v2

@@
 For example, in medical imaging, the primary objective is to obtain the most accurate reconstruction or best estimate of the underlying image, rather than to generate diverse samples from the posterior distribution.
-Optimization-based approaches—such as Resample (song2023solving, ), DiffPIR (zhu2023denoising, ), DCDP (li2024decoupled, ), and DMPlug (wang2024dmplug, )—have shown strong performance by alternating between denoising, optimization, and resampling to address inverse problems.
+Optimization-based approaches—such as Resample (Song et al., 2023c), DiffPIR (Zhu et al., 2023), DCDP (Li et al., 2024), and DMPlug (Wang et al., 2024)—have shown strong performance by alternating between denoising, optimization, and resampling to address inverse problems.
 Unlike DPS, which attempts to sample from the posterior distribution $p(x_{0}\mid y)$, optimization-based approaches prioritize reconstruction performance over distributional faithfulness.
@@
 We analyze its relationship to global MAP and DPS, and show that LMAPS unifies Tweedie Moment Projected Diffusion (TMPD) and optimization-based inverse problem methods under a single framework.
-The relationship between LMAPS and existing methods are presented in Figure 1.
+The relationship between LMAPS and existing methods are presented in fig.
+1.
   • Methodological.
@@
 However, the additional term $\nabla_{x_{t}}\log p(y\mid x_{t})$ is still intractable.
-One can train a neural network to approximate $\nabla_{x_{t}}\log p(y\mid x_{t})$, like classifier guidance (dhariwal2021diffusion, ).
-Training-free guidance, such as in (chung2022diffusion, ), usually approximates $\nabla_{x_{t}}\log p(y\mid x_{t})$ by a convenient single-sample approximation, $p(y\mid x_{t})\approx p(y\mid m_{0\mid t}(x_{t}))$, according to chain rule:
+One can train a neural network to approximate $\nabla_{x_{t}}\log p(y\mid x_{t})$, like classifier guidance (Dhariwal and Nichol, 2021).
+Training-free guidance, such as in (Chung et al., 2022), usually approximates $\nabla_{x_{t}}\log p(y\mid x_{t})$ by a convenient single-sample approximation, $p(y\mid x_{t})\approx p(y\mid m_{0\mid t}(x_{t}))$, according to chain rule:
 $\nabla_{x_{t}}\log p(y\mid x_{t})\approx\nabla_{x_{t}}m_{0\mid t}(t,x_{t})\nabla_{m_{0\mid t}}\log p(y\mid m_{0\mid t}(t,x_{t})).$ (5)
-Decoupled Annealing Posterior Sampling (DAPS) (zhang2025improving, ).
+Decoupled Annealing Posterior Sampling (DAPS) (Zhang et al., 2025a).
 Alternatively, DAPS developed a new framework to sample $x_{0}\sim\pi_{0\mid y}(x_{0}\mid y)$, which is given by the following iterations:
@@
 Local MAP equals DPS only in Gaussian conditional settings; outside them, local MAP generally does not sample the posterior and can fail to reach the global MAP.
-We visualize a toy example in Figure 3.
+We visualize a toy example in fig.
+3.
 Compared to DPS and DAPS, LMAPS is less likely to generate samples in between-mode regions or low-density regions.
@@
 The final ingredient for constructing a local posterior and solving the resulting MAP problem is the choice of prior $p(x_{0}\mid x_{t})$.
-While the true transition kernel of a diffusion model prior requires simulation, we can proceed as in previous work (boys2023tweedie, ; song2023pseudoinverse, ) by projecting onto the first two moments using a Gaussian approximation, $p(x_{0}\mid x_{t})\approx\mathcal{N}(x_{0};m_{0\mid t},\Sigma_{0\mid t})$, where $m_{0\mid t}(x_{t}):=\mathbb{E}[x_{0}\mid x_{t}]$.
-While boys2023tweedie show that $\Sigma_{0\mid t}^{\textsc{tmpd}}(x_{t}):=\mathbb{E}[(x_{0}-m_{0\mid t})(x_{0}-m_{0\mid t})^{T}\mid x_{t}]=\frac{\sigma_{t}^{2}}{\alpha_{t}}\nabla_{x_{t}}m_{0|t}$ to faithfully reflect the covariance of $x_{0}\mid x_{t}$, we will consider flexible choices of $\Sigma_{0|t}$.
+While the true transition kernel of a diffusion model prior requires simulation, we can proceed as in previous work (Boys et al., 2023; Song et al., 2023b) by projecting onto the first two moments using a Gaussian approximation, $p(x_{0}\mid x_{t})\approx\mathcal{N}(x_{0};m_{0\mid t},\Sigma_{0\mid t})$, where $m_{0\mid t}(x_{t}):=\mathbb{E}[x_{0}\mid x_{t}]$.
+While Boys et al.
+(2023) show that $\Sigma_{0\mid t}^{\textsc{tmpd}}(x_{t}):=\mathbb{E}[(x_{0}-m_{0\mid t})(x_{0}-m_{0\mid t})^{T}\mid x_{t}]=\frac{\sigma_{t}^{2}}{\alpha_{t}}\nabla_{x_{t}}m_{0|t}$ to faithfully reflect the covariance of $x_{0}\mid x_{t}$, we will consider flexible choices of $\Sigma_{0|t}$.
 Finally, the local MAP problem amounts to solving
@@
 In the implementation, we adopt gradient descent to solve $x_{0}^{*}$ in Eq. 13, the algorithm of LMAPS for inverse problems is provided in algorithm 4.
-Relationship to optimization-based methods. Previous optimization-based approaches (song2023solving, ; li2024decoupled, ; zhu2023denoising, ) solve for $x_{0}^{*}$ through the following objective:
+Relationship to optimization-based methods. Previous optimization-based approaches (Song et al., 2023c; Li et al., 2024; Zhu et al., 2023) solve for $x_{0}^{*}$ through the following objective:
 $x_{0}^{*}=\arg\min\|x_{0}-m_{0\mid t}\|^{2}+\lambda_{t}\|y-\mathcal{H}(x_{0})\|^{2},$ (14)
@@
 3.2, the local MAP solution matches the posterior mean for Gaussian posteriors $p(x_{t}\mid x_{0},y)$ arising from linear inverse problems $p(y\mid x_{0})=\mathcal{N}(Hx_{0},\sigma_{y}^{2}\mathbb{I})$ with a Gaussian assumption on the prior $p(x_{0}\mid x_{t})=\mathcal{N}(x_{0};m_{0\mid t},\Sigma_{0\mid t})$.
-Solving in closed form for the posterior mean as in (boys2023tweedie, ), we have
+Solving in closed form for the posterior mean as in (Boys et al., 2023), we have
 $x_{0}^{*}=m_{0\mid t}+\Sigma_{0\mid t}H^{T}(H\Sigma_{0\mid t}H^{T}+\sigma_{y}^{2}\mathbb{I})^{-1}(y-Hm_{0\mid t}).$ (15)
-We recover Tweedie Moment-Projected Diffusion (boys2023tweedie, ) as a special case for $\Sigma^{\textsc{tmpd}}_{0\mid t}=\frac{\sigma_{t}^{2}}{\alpha_{t}}\nabla_{x_{t}}m_{0\mid t}(x_{t})$, which is expensive since it requires the gradient with respect to the denoiser $m_{0\mid t}$.
+We recover Tweedie Moment-Projected Diffusion (Boys et al., 2023) as a special case for $\Sigma^{\textsc{tmpd}}_{0\mid t}=\frac{\sigma_{t}^{2}}{\alpha_{t}}\nabla_{x_{t}}m_{0\mid t}(x_{t})$, which is expensive since it requires the gradient with respect to the denoiser $m_{0\mid t}$.
 When applying LMAPS to linear inverse problems, we assume $\Sigma_{0|t}=\frac{k}{\text{SNR}_{t}}\mathbb{I}$ as in Sec.
@@
 For nonlinear image restoration, we consider (1) phase retrieval, (2) high dynamic range (HDR) reconstruction, (3) nonlinear deblurring, (4) JPEG restoration, (5) quantization, where HDR, JPEG restoration and quantization are nonlinear inverse problems with non-differentiable operators.
-For scientific inverse problems, we adopt the benchmark from InverseBench (zheng2025inversebench, ), which includes Linear Inverse Scattering (LIS), Compressed sensing MRI (CS-MRI) and Black Hole Imaging.
+For scientific inverse problems, we adopt the benchmark from InverseBench (Zheng et al., 2025), which includes Linear Inverse Scattering (LIS), Compressed sensing MRI (CS-MRI) and Black Hole Imaging.
 More details are provided in the App.
@@
 Dataset.
-For image restoration, we evaluated our method on FFHQ (karras2019style, ) $256\times 256$ and ImageNet $256\times 256$ datasets (deng2009imagenet, ).
+For image restoration, we evaluated our method on FFHQ (Karras et al., 2019) $256\times 256$ and ImageNet $256\times 256$ datasets (Deng et al., 2009).
 Following DAPS, we test the same subset of 100 images for both datasets.
-For scientific inverse problems, we adopt the same dataset as InverseBench (zheng2025inversebench, ).
+For scientific inverse problems, we adopt the same dataset as InverseBench (Zheng et al., 2025).
 Pretrained models.
-For image restoration tasks, we utilize the pre-trained checkpoint (chung2022diffusion, ) on the FFHQ dataset and the pre-trained checkpoint (dhariwal2021diffusion, ) on the ImageNet dataset.
+For image restoration tasks, we utilize the pre-trained checkpoint (Chung et al., 2022) on the FFHQ dataset and the pre-trained checkpoint (Dhariwal and Nichol, 2021) on the ImageNet dataset.
 For scientific inverse problems, we adopt the pre-trained checkpoints from InverseBench.
@@
 Metrics.
-For image restoration tasks, we report Peak Signal-to-Noise Ratio (PSNR), Structural SIMilarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) (zhang2018unreasonable, ).
+For image restoration tasks, we report Peak Signal-to-Noise Ratio (PSNR), Structural SIMilarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018).
 For scientific inverse problems, we primarily present PSNR in the main text, while additional task-specific metrics are provided in the App.
@@
 Ablation studies.
-Figure 4 presents ablation studies on optimization steps across different diffusion steps.
+fig.
+4 presents ablation studies on optimization steps across different diffusion steps.
 The best performance is typically observed at NFE = 200–500, where increasing the number of optimization steps per diffusion step.
@@
 6 Related work
-Recent advances in conditional generation have led to breakthroughs in text-to-image synthesis, semantic editing, and domain-specific applications such as image-to-image translation and controlled signal reconstruction (song2023loss, ; ye2024tfg, ; skreta2025feynman, ; singhal2025general, ; zheng2023guided, ).
-These methods have been especially impactful in solving inverse problems, including image restoration and scientific reconstruction tasks (wang2022zero, ; zheng2025inversebench, ).
-A wide range of approaches have been developed, spanning linear projection methods (wang2022zero, ; kawar2022denoising, ; zhang2025measurement, ; dou2024diffusion, ), Monte Carlo sampling (wu2023practical, ; phillips2024particle, ), variational inference (feng2023score, ; mardani2023variational, ; janati2024divide, ), and optimization-based strategies (song2023solving, ; zhu2023denoising, ; li2024decoupled, ; wang2024dmplug, ; alkhouri2024sitcom, ; he2023manifold, ).
-Among these, Diffusion Posterior Sampling (DPS) and its variants (zhang2025improving, ; chung2022diffusion, ; song2023loss, ; yu2023freedom, ; rout2024beyond, ; yang2024guidance, ; bansal2023universal, ; boys2023tweedie, ; song2023pseudoinverse, ; ho2022classifier, ) have gained wide adoption due to their strong empirical performance and interpretability, as they directly sample from the posterior distribution $p(x_{0}\mid y)$.
+Recent advances in conditional generation have led to breakthroughs in text-to-image synthesis, semantic editing, and domain-specific applications such as image-to-image translation and controlled signal reconstruction (Song et al., 2023a; Ye et al., 2024; Skreta et al., 2025; Singhal et al., 2025; Zheng et al., 2023).
+These methods have been especially impactful in solving inverse problems, including image restoration and scientific reconstruction tasks (Wang et al., 2022; Zheng et al., 2025).
+A wide range of approaches have been developed, spanning linear projection methods (Wang et al., 2022; Kawar et al., 2022; Zhang et al., 2025b; Dou and Song, 2024), Monte Carlo sampling (Wu et al., 2023; Phillips et al., 2024), variational inference (Feng et al., 2023; Mardani et al., 2023; Janati et al., 2024), and optimization-based strategies (Song et al., 2023c; Zhu et al., 2023; Li et al., 2024; Wang et al., 2024; Alkhouri et al., 2024; He et al., 2023).
+Among these, Diffusion Posterior Sampling (DPS) and its variants (Zhang et al., 2025a; Chung et al., 2022; Song et al., 2023a; Yu et al., 2023; Rout et al., 2024; Yang et al., 2024; Bansal et al., 2023; Boys et al., 2023; Song et al., 2023b; Ho and Salimans, 2022) have gained wide adoption due to their strong empirical performance and interpretability, as they directly sample from the posterior distribution $p(x_{0}\mid y)$.
 More recently, attention has shifted toward maximum a posteriori (MAP) estimation with diffusion priors.
-xu2025rethinking argued that DPS is in fact more consistent with the principle of MAP estimation rather than true posterior sampling, although their proposed sampling algorithm differs from ours.
-Finally, gutha2025inverse proposed sampling from the global MAP solution, $\arg\max p(x_{0}\mid y)$, though their approach is largely restricted to linear inverse problems.
+Xu et al.
+(2025) argued that DPS is in fact more consistent with the principle of MAP estimation rather than true posterior sampling, although their proposed sampling algorithm differs from ours.
+Finally, Gutha et al.
+(2025) proposed sampling from the global MAP solution, $\arg\max p(x_{0}\mid y)$, though their approach is largely restricted to linear inverse problems.
 7 Conclusion
@@
 C.1 Dataset details
-For scientific inverse problems, we adopt fluorescence microscopy images from InverseBench [40] on linear inverse scattering tasks, General Relativistic MagnetoHydroDynamic (GRMHD) [20] on black hole imaging, and multi-coil raw $k$-space data from the fastMRI knee dataset [36] on CS-MRI.
+For scientific inverse problems, we adopt fluorescence microscopy images from InverseBench [Zheng et al., 2025] on linear inverse scattering tasks, General Relativistic MagnetoHydroDynamic (GRMHD) [Mizuno, 2022] on black hole imaging, and multi-coil raw $k$-space data from the fastMRI knee dataset [Zbontar et al., 2018] on CS-MRI.
 C.2 Inverse problem details
-Baselines from DAPS [37].
+Baselines from DAPS [Zhang et al., 2025a].
 For image restoration tasks include: (1) super-resolution, (2) Gaussian deblurring, (3) motion deblurring, (4) inpainting (with a box mask), and (5) inpainting (with a 70% random mask), (6) phase retrieval, (7) high dynamic range (HDR) reconstruction, (8) nonlinear deblurring, we follow the same experimental setup as in DAPS.
-InverseBench [40]. For scientific inverse problems, we adopt the setting introduced in InverseBench.
+InverseBench [Zheng et al., 2025]. For scientific inverse problems, we adopt the setting introduced in InverseBench.
 JPEG Restoration. We address JPEG restoration with quality factors of $\mathrm{QF}=5$.
@@
 C.3 Baseline details
-For SITCOM [1], we use the hyperparameter configuration recommended in the original paper, with $N=20$ and $K=30$, resulting in $600$ NFEs and requiring gradient computation with respect to the U-Net.
+For SITCOM [Alkhouri et al., 2024], we use the hyperparameter configuration recommended in the original paper, with $N=20$ and $K=30$, resulting in $600$ NFEs and requiring gradient computation with respect to the U-Net.
 […]
-Other baselines we adopt the same reported results as in DAPS [37] and InverseBench [40].
+Other baselines we adopt the same reported results as in DAPS [Zhang et al., 2025a] and InverseBench [Zheng et al., 2025].
 C.4 Complete List of hyper-parameters